OpenStudy (anonymous):

What is the graph of 5x=2y-20 ?

sam (.sam.):

First arrange the equation in the form y=mx+c

OpenStudy (anonymous):

y=-10-5x over 2

sam (.sam.):

5x=2y-20 add 20 to both sides \[5x+20=2y\] divide 2 both sides \[y=\frac{5}{2}x+10\]

OpenStudy (anonymous):

ooooooo ok

OpenStudy (anonymous):

hold real quick please

OpenStudy (anonymous):

will that be my graph ?

sam (.sam.):

yes

OpenStudy (anonymous):

r sure

sam (.sam.):

you can verify it by using some coordinates, try using x=-2 , x=-1, x=0, x=1, x=2 on the equation. When 'x' has these points, find the 'y' values by substituting them into the equation and then you'll get the coordinates to plot.

OpenStudy (anonymous):

ok thankssssssssss

sam (.sam.):

welcome